OpenStudy (sonyalee77):

To divide a decimal by a power of ten, move the decimal point to the right the same number of places as there are zeros in the power of ten. A. True B. False

OpenStudy (candy13106):



OpenStudy (anonymous):

10^-1 = .1 = 1/10 10.0 * 10^-1 = 1.0 1.0 *10^-1 = 0.1

OpenStudy (anonymous):

Here's a trick to make it easy for you: Divide by ten = left Multiply by ten = right

OpenStudy (candy13106):